Fred needs 55 ml of 24% sucrose solution to do his science experiment. He needs to mix a 15% sucrose solution with some 40% sucrose solution to get the 24% solution he wants. How much of each should Fred use? (Include units with your numerical answers.)
Answer by richwmiller(17219)

a+b=55,
0.15*a+0.4*b=0.24*55
a=55-b
0.15*(55-b)+0.4*b=13.2
8.25-0.15b+0.4*b=13.2
0.25*b=4.95
b=19.8
a=55-b
a=35.2 mL at 15%
b=19.8 mL at 40%
check
0.15*35.2+0.4*19.8=0.24*55
5.28+7.92=13.2
13.2=13.2
ok